# Enhancing Android Malware Detection: The Influence of ChatGPT on Decision-centric Task

### v1

@@ 7. Related work
Android malware detection: Android malware detection techniques can be categorized into two forms: static analysis and dynamic analysis. Static analysis involves decompiling the APK file to obtain the source code and metadata files, followed by extraction of significant features for malware detection and analysis such as the application programming interface (Liu et al., 2021a; Tao et al., 2018), permissions (L. et al., 2021; Dharmalingam and Palanisamy, 2020; Zhi et al., 2018; Alazab et al., 2020), intent (Taheri et al., 2020), function call graph (Fan et al., 2018; Liu et al., 2021b; Gao et al., 2021), control flow (Ma et al., 2019), and grayscale (Bakour and Nver, 2021). Dynamic analysis (Bae and Shin, 2016; Saracino et al., 2018; Vp et al., 2019; Zhou et al., 2019; Surendran et al., 2020; Massarelli et al., 2020) involves the extraction of features by monitoring the application’s execution in real-time. With the development of artificial intelligence, many related researches focus on models, such as SVM (Liu et al., 2021b; Varna and Visalakshi, 2020; Salah et al., 2020), and decision tree (DT) (Nisa et al., 2020), RF (L. et al., 2021; Zhu et al., 2017), AdaBoost (Bakour and Nver, 2021) and ensemble learning algorithm (Yerima et al., 2019; Feng et al., 2018), to classify malware. In addition, even deep learning methods (Zhu et al., 2021; Au et al., 2012) and meta-learning methods (Li et al., 2022a; Zhu et al., 2022) are directly used for classification. Although these methods demonstrate superior capabilities, they lack the ability to interpret and analyze.
@@ 8. Conclusion and Future Work
ChatGPT has demonstrated exceptional interpretation and analysis capabilities, making it applicable to a variety of tasks. We are the first to study the influence of ChatGPT on Android malware detection tasks. We select state-of-the-art Android malware detection solutions for comparison with ChatGPT. The research indicates that existing detection solutions suffer not only from dataset bias but also lack explanation and analysis capabilities. Given the existence of these problems and the excellent interpretability of ChatGPT, we identify a novel perspective for detecting Android malware, that is, less decision-making, and more explanation. In the future, we plan to design a specialized large language model specifically for Android malware detection to enhance interpretability.

### v2

@@ 6.1. Analysis
6.2. Why ChaGPT
We select ChatGPT as the LLM for the following reasons.
First, the goal of our research is not to compare or determine which LLM is superior in the field of Android malware detection and analysis. Instead, our focus is to demonstrate that LLMs have the potential to transform the existing detection paradigm. We aim to inform developers and users that malware detection methodologies in the LLM era should evolve. Rather than solely relying on traditional decision-oriented methods, greater emphasis should be placed on providing analysis and explanations to enhance understanding and usability.
Second, This research began in 2023. At the time, Claude 1.0 was primarily designed as a chatbot with limited functionality. When Claude 2.0 was released, its performance fell short of ChatGPT 4 (Borji and Mohammadian, 2023). It was not until June 2024, with the release of Claude 3.5 Sonnet, that its performance surpassed ChatGPT 4  [^7] https://www.anthropic.com/news/claude-3-5-sonnet. However, Claude imposed strict usage restrictions, including requirements for billing and IP address consistency, as well as limitations on access in certain regions. Since our primary authors are based in Greater China, these restrictions made it impractical to utilize Claude in our research. In contrast, ChatGPT’s more lenient access policies allowed us to integrate it seamlessly into our workflow.
Finally, due to hardware constraints and the lack of comprehensive datasets (not only in terms of sample size but also missing contextual information such as application background data, etc.), training a large-scale open-source LLM was not feasible. ChatGPT offered a ready-to-use solution with robust capabilities, enabling us to conduct our research effectively. That said, we are actively collecting relevant datasets to support the future development of an open-source LLM specifically designed for Android malware detection.
These considerations made ChatGPT the most convenient and accessible option for our study. Looking ahead, we plan to build upon this work and develop a specialized LLM for Android malware detection using open-source models.
@@ 7. Related work
Android malware detection: Android malware detection techniques can be categorized into two forms: static analysis and dynamic analysis. Static analysis involves decompiling the APK file to obtain the source code and metadata files, followed by extraction of significant features for malware detection and analysis such as the application programming interface (Liu et al., 2021a; Tao et al., 2018), permissions (L. et al., 2021; Dharmalingam and Palanisamy, 2020; Zhi et al., 2018; Alazab et al., 2020), intent (Taheri et al., 2020), function call graph (Fan et al., 2018; Liu et al., 2021b; Gao et al., 2021), control flow (Ma et al., 2019), and grayscale (Bakour and Nver, 2021). Li et al. (Li et al., 2024) propose a meta-learning-based multi-classifier that uses multiple features such as permissions and APIs to classify malware and achieves good results. Although the meta-classifier learns meta-knowledge in the dataset, it cannot provide a basis for making decisions. Dynamic analysis (Bae and Shin, 2016; Saracino et al., 2018; Vp et al., 2019; Zhou et al., 2019; Surendran et al., 2020; Massarelli et al., 2020) involves the extraction of features by monitoring the application’s execution in real-time. Ficco (Ficco, 2022) combines general and specialized detectors to enhance the randomness of detection and improve the overall detection rate. Although this solution combines static and dynamic analysis, it is still decision-oriented and does not consider the interpretability of the solution. With the development of artificial intelligence, many related researches focus on models, such as SVM (Liu et al., 2021b; Varna and Visalakshi, 2020; Salah et al., 2020), and decision tree (DT) (Nisa et al., 2020), RF (L. et al., 2021; Zhu et al., 2017), AdaBoost (Bakour and Nver, 2021) and ensemble learning algorithm (Yerima et al., 2019; Feng et al., 2018), to classify malware. In addition, even deep learning methods (Zhu et al., 2021; Au et al., 2012) and meta-learning methods (Li et al., 2022a; Zhu et al., 2022) are directly used for classification. Although these methods have shown superior capabilities in detecting malware, they lack the ability to explain and analyze, and are unable to analyze the features of each input and inform developers and users of the functions and potential threats represented by these features.
@@ 8. Conclusion and Future Work
ChatGPT has demonstrated exceptional interpretation and analysis capabilities, making it applicable to a variety of tasks. We are the first to study the influence of ChatGPT on Android malware detection tasks. We select state-of-the-art Android malware detection solutions for comparison with ChatGPT. The research indicates that existing detection solutions suffer not only from dataset bias but also lack explanation and analysis capabilities. Given the existence of these problems and the excellent interpretability of ChatGPT, we identify a novel perspective for detecting Android malware, that is, less decision-making, and more explanation. In the future, we aim to enhance and expand the existing datasets by addressing gaps and completing missing information, such as application background details. Additionally, we plan to develop a specialized large language model tailored for Android malware detection, with a focus on improving both detection accuracy and interpretability.